Shift summary for eve 9/24/99

Jason Hamm (jhamm@physics.arizona.edu)
Fri, 24 Sep 1999 21:58:36 -0700 (MST)


Shift Summary for eve shift, 09/24/99

On Shift: A. Erwin, J. Hamm

Hours of beam available at all intensities: 1.5
Hours of beam available at nominal intensity: 1.25
Hours of E832 data (special runs not included): ~1.25
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Recovering from MI power trip, linac problems,
problems with new memory module
========================================================================
Runs ended this shift:

14713 - Beam_799

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Very little data taking, with some memory problems to
keep things a little exciting.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 MI still recovering from earlier power fault.

20:45 Beam arrives, starting at 4e12, but working up to
8e12 over the next 10 minutes.

20:50 Onsplit requests UPPE01 on ktev3.

~21:30 Notice murmur window is flooded with ktevana
errors at start of each spill. All errors are
in stream 5, plane 4. The raw data looks
corrupted. l3#stat says there are serious
errors for plane 4.

Furthermore, both the murmur and the l3#stat
errors exist in runs going back through last
night, ever since the new memories were added.
We conclude that one of these memories is
probably bad, and page JRL.

Stop run 14713.

Beam is gone for 20 minutes while JRL replaces
the first memory module for stream 5, plane 4.

22:15 Start Beam_799 run 14714 after memory swap -
problem seems to be fixed.

22:40 Beam gone, linac problems - est 1 hr

23:10 Beam back

23:30 Beam gone, linac problems - est 20 min

23:45 Beam returns, but very intermittent through
end of shift.

23:55 Hogan raises laser intensity by 2000 ticks.

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

The most probable reason for the memory problems not being caught last
night is that ktevana audio alarms are only issued for errors on the
monitor plane, and not the filter plane. This should be looked into.